OpenStudy (anonymous):

Can someone explain to me how the compound interest formula works? (C=(p+1)^n)

OpenStudy (irishboy123):

Starting capital X Interest per period p After 1 period you will have capital at start of period plus interest on capital = X + (X*p) = X(1+p) After next period you will have capital at start of period plus interest on capital = X(1+p) + X(1+p)* p = X(1+p)(1+p) And so on. So it's just a geometric series starting value X and with ratio 1+ p